A polygon has the sum of interior angles as 1440degree. Write the name of the polygon.
Answer by Cromlix(4381)

Formula = (180(n - 2)) = total interior angles
Where 'n' = no. of sides
(180(n - 2)) = 1440
180n - 360 = 1440

180n = 1440 + 360

180n = 1800
n = 10
Therefore your polygon is a decagon.
Hope this helps.
:-)